OpenStudy (nuttyliaczar):

Can anyone help me translate the equation for an electromagnetic wave into its parts?

OpenStudy (nuttyliaczar):

I'll be back in a bit but let me know if you're adept at reading these kinds of equations

OpenStudy (nuttyliaczar):

Excuse the horrific file name ^^'

OpenStudy (captainfluffy):

lol

OpenStudy (captainfluffy):

damn ok ill try my best

OpenStudy (captainfluffy):

i dont think i will be of any help.... damn... i... dont even know were to start, my math skills are useless

OpenStudy (captainfluffy):

is there anything else maybe?

OpenStudy (nuttyliaczar):

That's what I'm stuck on atm, so no just this

OpenStudy (captainfluffy):

man if only i had your knowledge i could try and help but i take basic math .

OpenStudy (nuttyliaczar):

If I was good at it I'd have done it by now =/

OpenStudy (unklerhaukus):

can you determine the angular frequency \(\omega\)?

OpenStudy (unklerhaukus):

for a wave oscillation with the angular frequency \(\omega\), one of the components in the complex exponential (or trig. function), must be \(\omega t\)

OpenStudy (nuttyliaczar):

Is it 9.42 X 10^15?

OpenStudy (unklerhaukus):

and with SI units?

OpenStudy (nuttyliaczar):

Radians per second?

OpenStudy (unklerhaukus):

yep

OpenStudy (unklerhaukus):

the other term in the complex exponential (or trig. function), is k•r, where k is the wave vector, and r is position vector

OpenStudy (nuttyliaczar):

Hey I do still need help with this but I'm extremely tired. Will you be around tomorrow?

OpenStudy (irishboy123):

👀

OpenStudy (johnweldon1993):

Oh fun haven't worked with these in a bit! :P The direction of oscillation will act in the direction of the amplitude (normalized) \[\large \frac{-6 \hat{i} + 3\sqrt{5}\hat{j}}{|(-6)^2 + (3\sqrt{5})^2|}\] Scalar Amplitude would just be \[\large |A| = (10^4 \frac{V}{m})\sqrt{(-6)^2 + (3\sqrt{5})^2}\] The direction of propagation works with the complex exponent, we know it is in the form \(\large e^{i[\vec{k} \cdot \vec{r}-\omega t]}\) and we know the position vector \(\large \vec{r} = x\hat{x} + y\hat{\hat{y}}\) we can see that \(\large \vec{k} = \frac{1}{3}\pi \times 10^7(\sqrt{5}\hat{i} + 2\hat{j})\) The propagation number 'k' is \(\large k = \sqrt{\vec{k} \cdot \vec{k}}\) which here means \(\large \sqrt{(\pi \times 10^7) \times (\pi \times 10^7)}\) or just \(\large \pi \times 10^7\) We know the wavelength can be found by \(\large \lambda = \frac{2\pi}{k}\) so I'm not gonna help you with that :P You already found the angular frequency to be 9.42 x 10^15 So we can use that to solve for the frequency since we know \(\large f = \frac{\omega}{2\pi}\) And finally the speed would be the product of the wavelength and the frequency \(\large v = \lambda f\)